# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v3

@@ 5.1 Entity Alignment in Knowledge Graphs
The entity alignment (EA) task aims to discover one-to-one equivalent entity pairs across two KGs that refer to the same real-world identity. Early EA models are probabilistic methods that compute similarities and perform equivalence reasoning in the input space. For example, PARIS [37] is an unsupervised ontology alignment model that jointly aligns entities, relations, and classes across KGs. It converts each relation into a logic-rule based function, and iteratively refines entity alignment probabilities using the functional and inverse functional properties of relations. PARIS is later extended to a semi-supervised EA model, PARIS+ [17], allowing the incorporation of prior seed alignments.
@@
Since 2017, most EA models are embedding-based, using distances between entity embeddings in latent spaces to measure the semantic correspondences between entities. Inspired by TransE [4], MTransE [7] embeds two KGs into two respective embedding spaces, where a transformation matrix is learned using prior seed alignments. To obtain better KG embeddings, TransEdge [41] enhances the translational scoring function by replacing the relation embedding with an edge embedding that incorporates information from both the head and tail entities, in addition to the relation information. To reduce the number of parameters involved, most subsequent models [39, 40, 62] embed KGs into a common latent space by imposing the embeddings of pre-aligned entities to be as close as possible. This ensures that alignment similarities between entities can be directly measured via their embeddings.
@@
More recent EA models leverage GNNs to incorporate KG structural information for entity alignment. For example, GCN-Align [52] adopts GCNs to learn better entity embeddings for alignment inference. However, GCNs and their variants are inclined to result in alignment conflicts, as their feature aggregation scheme incurs an over-smoothing issue [24, 13]: The embeddings of entities among local neighborhood become indistinguishable similar as the number of GCN layers increases. To mitigate the over-smoothing effect, more recent works [54, 53, 63] adopt a highway strategy [35] on GCN layers, which “mixes” the learned entity embeddings with the original features. Another line of research efforts is devoted to improving GCN-based approaches through considering heterogeneous relations in KGs. HGCN [54] jointly learns the embeddings of entities and relations, without considering the directions of relations. RDGCN [53] performs embedding learning on a dual relation graph, but fails to incorporate statistical information of neighboring relations of an entity. RNM [63] uses iterative relational neighborhood matching to refine finalized entity embedding distances. This matching mechanism proves to be empirically effective, but it is used only after the completion of model training and fails to reinforce embedding learning in turn. BERT-INT [44] leverages the BERT (Bidirectional Encoder Representations from Transformers) model to capture both entity and contextual information from relational paths between entities, thereby enhancing entity alignment performance by incorporating rich semantics such as entity descriptions. Yet, obtaining powerful description information in practice can be challenging in many real-world scenarios. All the aforementioned models, however, require an abundance of prior seed alignments provided for training purposes, which are labor-intensive and costly to acquire in real-world KGs.
@@
To tackle the shortage of prior seed alignments, semi-supervised EA models have been proposed in recent years. As a prominent learning paradigm among such, pseudo-labeling-based methods, e.g., BootEA [40], IPTransE [62], TransEdge [41], RNM [63], MRAEA [21], and CPL-OT [10], propose to iteratively pseudo-label unaligned entity pairs and add them to seed alignments for subsequent model training. For RNM, there is a slight difference that it augments seed alignments to rectify embedding distance after the completion of model training. Although these methods have achieved promising performance gains, the confirmation bias associated with iterative pseudo-labeling has been largely under-explored. Recent methods like RNM [63] and MRAEA [21] use simple heuristics to preserve only the most convincing alignment pairs, for example, those with the smallest distance, at the presence of conflicts. BootEA [40] and CPL-OT [10], on the other hand, model the inference of pseudo-labeled alignments as an assignment problem, where the most likely aligned pairs are selected at each pseudo-labeling iteration. Unlike BootEA that selects a small set of pseudo-labeled alignments using a pre-specified threshold, CPL-OT imposes a full match between two unaligned entity sets to maximize the number of pseudo-labeled alignments at each iteration. Both methods impose constraints to enforce hard alignments to alleviate alignment conflicts, which may potentially increase one-to-one misalignments.
@@ 5.3 Optimal Transport on Graphs
Optimal Transport (OT) is the general problem of finding an optimal plan to move one distribution of mass to another with the minimal cost [48]. As an effective metric to define the distance between probability distributions, OT has been applied in computer vision and natural language processing over a range of tasks including machine translation, text summarization, and image captioning [47, 6]. In recent years, OT has also been studied on graphs to match graphs with similar structures or align nodes/entities across graphs. For graph partitioning and matching, the transport on the edges across graphs is used to define the Gromov-Wasserstein (GW) discrepancy [46] that measures how edges in a graph compare to those in another graph [56, 23, 55]. For entity alignment across graphs, Pei et al [28] incorporate an OT objective into the overall loss to enhance the learning of entity embeddings. Tang et al [43] propose to jointly perform structure learning and OT alignment through minimizing multi-view GW distance matrices between two attributed graphs. These methods have primarily used OT to define a learning objective, which involves bi-level optimization for model training. To further enhance the scalability of OT modeling for entity alignment, Mao et al [22] propose to make the similarity matrix sparse by dropping its entries close to zero. However, this sparse OT modeling potentially violates the constraints of the OT objective, failing to guarantee one-to-one correspondences across two KGs. In our work, we focus on tackling the scarcity of prior seed alignments via iterative pseudo-labeling; we seek to find more accurate one-to-one alignment configurations between entities via OT modeling, thus eliminating conflicted misalignments at each pseudo-labeling iteration and mitigating confirmation bias.

### v4

@@ 5.1 Entity Alignment in Knowledge Graphs
The entity alignment (EA) task aims to discover one-to-one equivalent entity pairs across two KGs that refer to the same real-world identity. Early EA models are probabilistic methods that compute similarities and perform equivalence reasoning in the input space. For example, PARIS [37] is an unsupervised ontology alignment model that jointly aligns entities, relations, and classes across KGs. It converts each relation into a logic-rule based function, and iteratively refines entity alignment probabilities using the functional and inverse functional properties of relations. PARIS is later extended to a semi-supervised EA model, PARIS+ [17], allowing the incorporation of seed alignments.
@@
Since 2017, most EA models are embedding-based, using distances between entity embeddings in latent spaces to measure the semantic correspondences between entities. Inspired by TransE [4], MTransE [7] embeds two KGs into two respective embedding spaces, where a transformation matrix is learned using seed alignments. To obtain better KG embeddings, TransEdge [41] enhances the translational scoring function by replacing the relation embedding with an edge embedding that incorporates information from both the head and tail entities, in addition to the relation information. To reduce the number of parameters involved, most subsequent models [39, 40, 62] embed KGs into a common latent space by imposing the embeddings of pre-aligned entities to be as close as possible. This ensures that alignment similarities between entities can be directly measured via their embeddings.
@@
More recent EA models leverage GNNs to incorporate KG structural information for entity alignment. For example, GCN-Align [52] adopts GCNs to learn better entity embeddings for alignment inference. However, GCNs and their variants are inclined to result in alignment conflicts, as their feature aggregation scheme incurs an over-smoothing issue [24, 13]: The embeddings of entities among local neighborhood become indistinguishable similar as the number of GCN layers increases. To mitigate the over-smoothing effect, more recent works [54, 53, 63] adopt a highway strategy [35] on GCN layers, which “mixes” the learned entity embeddings with the original features. Another line of research efforts is devoted to improving GCN-based approaches through considering heterogeneous relations in KGs. HGCN [54] jointly learns the embeddings of entities and relations, without considering the directions of relations. RDGCN [53] performs embedding learning on a dual relation graph, but fails to incorporate statistical information of neighboring relations of an entity. RNM [63] uses iterative relational neighborhood matching to refine finalized entity embedding distances. This matching mechanism proves to be empirically effective, but it is used only after the completion of model training and fails to reinforce embedding learning in turn. BERT-INT [44] leverages the BERT (Bidirectional Encoder Representations from Transformers) model to capture both entity and contextual information from relational paths between entities, thereby enhancing entity alignment performance by incorporating rich semantics such as entity descriptions. Yet, obtaining powerful description information in practice can be challenging in many real-world scenarios. All the aforementioned models, however, require an abundance of seed alignments provided for training purposes, which are labor-intensive and costly to acquire in real-world KGs.
@@
To tackle the shortage of seed alignments, semi-supervised EA models have been proposed in recent years. As a prominent learning paradigm among such, pseudo-labeling-based methods, e.g., BootEA [40], IPTransE [62], TransEdge [41], RNM [63], MRAEA [21], and CPL-OT [10], propose to iteratively pseudo-label unaligned entity pairs and add them to seed alignments for subsequent model training. For RNM, there is a slight difference that it augments seed alignments to rectify embedding distance after the completion of model training. Although these methods have achieved promising performance gains, the confirmation bias associated with iterative pseudo-labeling has been largely under-explored. Recent methods like RNM [63] and MRAEA [21] use simple heuristics to preserve only the most convincing alignment pairs, for example, those with the smallest distance, at the presence of conflicts. BootEA [40] and CPL-OT [10], on the other hand, model the inference of pseudo-labeled alignments as an assignment problem, where the most likely aligned pairs are selected at each pseudo-labeling iteration. Unlike BootEA that selects a small set of pseudo-labeled alignments using a pre-specified threshold, CPL-OT imposes a full match between two unaligned entity sets to maximize the number of pseudo-labeled alignments at each iteration. Both methods impose constraints to enforce hard alignments to alleviate alignment conflicts, which may potentially increase one-to-one misalignments.
@@ 5.3 Optimal Transport on Graphs
Optimal Transport (OT) is the general problem of finding an optimal plan to move one distribution of mass to another with the minimal cost [48]. As an effective metric to define the distance between probability distributions, OT has been applied in computer vision and natural language processing over a range of tasks including machine translation, text summarization, and image captioning [47, 6]. In recent years, OT has also been studied on graphs to match graphs with similar structures or align nodes/entities across graphs. For graph partitioning and matching, the transport on the edges across graphs is used to define the Gromov-Wasserstein (GW) discrepancy [46] that measures how edges in a graph compare to those in another graph [56, 23, 55]. For entity alignment across graphs, Pei et al [28] incorporate an OT objective into the overall loss to enhance the learning of entity embeddings. Tang et al [43] propose to jointly perform structure learning and OT alignment through minimizing multi-view GW distance matrices between two attributed graphs. These methods have primarily used OT to define a learning objective, which involves bi-level optimization for model training. To further enhance the scalability of OT modeling for entity alignment, Mao et al [22] propose to make the similarity matrix sparse by dropping its entries close to zero. However, this sparse OT modeling potentially violates the constraints of the OT objective, failing to guarantee one-to-one correspondences across two KGs. In our work, we focus on tackling the scarcity of seed alignments via iterative pseudo-labeling; we seek to find more accurate one-to-one alignment configurations between entities via OT modeling, thus eliminating conflicted misalignments at each pseudo-labeling iteration and mitigating confirmation bias.